OpenStudy (anonymous):

Can anyone find the coordinates of the vertex of the graph of y=2x^2+4x-2

jimthompson5910 (jim_thompson5910):

y=2x^2+4x-2 is in the form y=ax^2+bx+c where a = 2 b = 4 c = -2

jimthompson5910 (jim_thompson5910):

Plug a = 2 and b = 4 into the formula x = -b/(2a) to get the x coordinate of the vertex

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (jdoe0001):

\(\bf \textit{vertex of a parabola}\\ \quad \\ y = {\color{red}{ 2}}x^2{\color{blue}{ +4}}x{\color{green}{ -2}}\qquad\quad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\) as jim_thompson5910 suggested

OpenStudy (anonymous):

I think it equals -1 Jim Thompson

OpenStudy (anonymous):

Is that the answer?

jimthompson5910 (jim_thompson5910):

x = -1 is the x coordinate of the vertex that's half of your answer

jimthompson5910 (jim_thompson5910):

you need the y coordinate your answer will be of the form (x,y)

OpenStudy (anonymous):

So is it (0,-1)

jimthompson5910 (jim_thompson5910):

plug x = -1 into y=2x^2+4x-2

OpenStudy (anonymous):

Oh so y=-4

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Or wrong :(

OpenStudy (jdoe0001):

ahemm

jimthompson5910 (jim_thompson5910):

y=2x^2+4x-2 y=2(-1)^2+4(-1)-2 y = -4 so you are correct

jimthompson5910 (jim_thompson5910):

the vertex is therefore the point (-1,-4)

OpenStudy (anonymous):

Thanks Jim I really appreciate it and I also thank you too Jdode

jimthompson5910 (jim_thompson5910):

you're welcome